# thanks: These three authors contributed equallythanks: These three authors contributed equallythanks: These three authors contributed equally

### v1

@@ IV Results and discussion
Let us now compare the magnetic-field dependence of the $\alpha$ mode energy with the proposed theoretical model. The observed blueshift, approximately quadratic in $B$, matches well the expectations raised by Eq. 3 in the limit of a relatively weak magnetic field. The validity of the theoretical model can be further corroborated using an additional set of data collected at various angles $\theta_{B}$ between $\mathbf{B}$ and the $c$ axis. Theoretically, we expect a weakening of the blueshift with increasing $\theta_{B}$, to gradually approach Eq. 5 with a nearly vanishing $B$-dependence for $E_{\mathrm{an}}\ll E_{J}$. Experimentally, we have indeed found that the blueshift of the $\alpha$ mode gradually flattens with increasing $\theta_{B}$. We illustrate this in Fig. 4e on the set of data collected at $\theta_{B}=50^{\circ}$. The complete set of data collected at various angles $\theta_{B}$ is presented in the Appendix C.
@@
To account for this, we have fitted the experimentally observed blueshift, see the dashed line in Fig. 4c, using the formula $(E_{\alpha}^{2}+\xi^{2}\mu_{B}^{2}B^{2})^{1/2}$, cf. Eq. 3. The extracted parameter $\xi=1.99\pm 0.03$ then allows us to calculate an approximate effective relation between the magnetic anisotropy and the factor $g$: $D\approx(\sqrt{3}E_{\alpha}/15)\sqrt{1-\xi^{2}/g^{2}}$. It is also worth noting that the parameter $\xi$ corresponds to the lowest possible value of the $g$ factor consistent with our magneto-optical data. SOC in Mn-based compounds is typically weak and the effective $g$ factor should not significantly deviate from 2. Setting the maximum value at $g=2.1$, we get the following upper limit for the out-of-plane magnetic anisotropy, $D<150$ $\mu$eV.
@@
Another independent estimate of magnetic anisotropy can be obtained thanks to the AFMR experiment in the SF configuration (${\mathbf{B}}\perp c$). There, the $\alpha$ mode is expected to redshift approximately by $2(g\mu_{B}BD)^{2}/E^{3}_{\alpha}$ as one can infer from Eq. 5. Since no such shift was observed experimentally, we conclude it to be smaller than our experimental resolution (0.5 cm-1 @ 16 T). In this way, we obtain the upper bound for the magnetic anisotropy: $D<250$ $\mu$eV. This estimate is less strict, but consistent with the one based on our AFMR experiments in the EC configuration.
@@
Using exchange parameters available in the literature for $\alpha$-MnTe we arrive at another way to obtain an estimate of the magnetic anisotropy. Szuszkiewicz et al. [25] mapped the magnon dispersion using inelastic neutron scattering and compared their results with the dispersion obtained by the LSW theory, considering a spin Hamiltonian equivalent to Eq. (1). They found good agreement with the theory, except for small departures around zone boundaries, and gave solid estimates of the corresponding exchange coupling constants. However, magnetic anisotropy seems to be the least precise parameter deduced, mainly due to the limited data set around $k=0$. Combining the exchange parameters $J_{1}$ and $J_{3}$ from Szuszkiewicz et al. [25] leads to $E_{J}\approx 33$ meV (this is essentially the magnon bandwidth seen in Fig. 5 of that reference) and using the $\alpha$ mode energy deduced from the optical measurements in this work (see Fig. 2), we conclude that the out-of-plane component of the magnetic anisotropy falls within $D=40\pm 10$ $\mu$eV [$E_{\mathrm{an}}=100\pm 25$ $\mu$eV]. This, in turn, allows us to estimate the effective $g$ factor, $g=2.03\pm 0.01$, using the AFMR data collected in the EC configuration, see the discussion above.
@@ V Conclusions
A long-wavelength ($k=0$) magnon mode was identified in the THz response of bulk MnTe using the experimentally observed dependence of its energy on the applied magnetic field and temperature. The observed behavior aligns well with predictions from a simplified model of an easy-plane (hard-axis) antiferromagnet, developed within the framework of linear spin-wave theory. Its energy at low temperatures, 3.5 meV, ranks MnTe among antiferromagnetic materials with moderate magnon gap (smaller than in some uniaxial systems [48, 49] but still appreciable). On the other hand, the magnon remains observable even at room temperature which is close to $T_{N}$ of hexagonal MnTe. Combining our experimental results with the values of the exchange parameters published earlier, we determined the out-of-plane component of the magnetic anisotropy as $D=40\pm 10$ $\mu$eV and the effective $g$ factor as $g=2.03\pm 0.01$.
@@ Acknowledgements.
We acknowledge discussions with M. Potemski. The work was supported by the Czech Science Foundation, project No. 23-04746S and TERAFIT project No. CZ.02.01.01/00/22__\__008/0004594 funded by OP JAK, call Excellent Research. Additional magnetometry on our bulk samples was performed by O. Kaman. J.D. and M.O. acknowledge the support received through the ANR-22-EXES-0001 (project CEQAS). Authors acknowledge the support of the LNCMI-CNRS, a member of the European Magnetic Field Laboratory (EMFL). Supplemental experiments were carried out in MGML (mgml.eu), supported by the Ministry of Education, Youth and Sports, the Czech Republic within the program of the Czech Research Infrastructures (project No. LM2023065). P.K. and L.N. ackowledge funding from the Grant Agency of the Charles University (grants No. 166123 and SVV--2024--260720).
@@ Appendix A Linear spin wave theory
The classical ground state configuration has to be determined first. Writing the classical energy ${\cal E}$ associated with the Hamiltonian (1) we get for $\mathcal{E}/S$
@@
where $\mathbf{m}_{i}$ denote unit vectors corresponding to the respective sublattice magnetizations. The external magnetic field is $\mathbf{B}=\mathbf{b}B$, and without loss of generality, the unit vector $\mathbf{b}$ is taken to be in the plane $\{\mathbf{x},\mathbf{z}\}$. The quantities $E_{\mathrm{an}}$, $E_{Z}$ and $E_{J}$ are defined in the main text.
@@ Appendix B Temperature dependence of electrical resistivity
The temperature dependence of the electrical resistivity for samples A and B is plotted in Fig. 5. Qualitatively, both samples show the same behaviour with increasing $T$. A sharp decrease in resistance, with the minima below 100 K, and a gradual increase at higher temperatures. At high temperatures, the resistance of sample B appears to be approximately one order of magnitude higher than that of sample A. Assuming that the scattering of free carriers at high temperatures is dominated by intrinsic mechanisms (optical phonons), the difference in the resistance values may suggest a higher carrier density in sample A.
@@
Figure 5: Resistivity dependence on temperature for samples A and B with carrier freeze-out evidenced at low temperatures. The experiments were carried out using the conventional lock-in technique at low frequencies. The profile of $R(T)$ curve qualitatively match overall transmission of MnTe samples in the THz range, as discussed in the main text.
Appendix C Magneto-transmission of $\alpha$-MnTe –
a complementary set of data
An additional set of magneto-transmission data collected in the Faraday configuration on sample A using the frequency-domain spectroscopy, with the magnetic field applied at different angles $\theta_{B}$ with respect to the $c$ axis of MnTe, is plotted in Fig. 6. This set of data illustrates the relatively weak sensitivity of the $\alpha$ mode to the in-plane component of the applied magnetic field, in line with the expectations of the theoretical model (Sect. II and Appendix A). The blue shift of the $\alpha$ mode induced by the magnetic field, well-pronounced at low angles $\theta_{B}$, becomes at the limit of our experimental resolution above $\theta_{B}=60^{\circ}$. At higher angles, no blueshift was observed, including the limiting case of $\theta_{B}=90^{\circ}$ which we probed in the Voigt configuration (data not presented).
Figure 6: Relative magneto-transmission, $T_{B}/T_{B=0}$ collected in the Faraday configuration on sample A at liquid helium temperature, plotted as color plots for several values of the angle between the magnetic field and the $c$ axis of MnTe: $\theta_{B}=0,20,35,50,60$ and 70 in panels (a) to (f). The data in panels (a) and (d) are the same as in Figs. 4a and e, but this time normalized by the zero-field transmission. In this way, the dark and light colors correspond to areas of $B$-induced absorption and transmission, respectively.
Appendix D Description of experimental techniques
@@
Fourier transform magneto-spectroscopy: Nonpolarized radiation from mercury lamp was analyzed by the Bruker Vertex 80v spectrometer and delivered, via light-pipe optics, to the sample which was kept at $T=4.2$ K in the helium exchange gas and placed inside a cryostat in a superconducting or resistive coil, below and above 16 T, respectively. In the superconducting coil, the radiation was then detected by a composite bolometer placed just below the sample. In this case, both Faraday and Voigt configurations were possible, with the light propagating along or perpendicular to the applied magnetic field, respectively. In the resistive coil, measurements were made only in the Faraday configuration. In order to reduce the noise level induced by the cooling water, the sample was placed on a mirror and the radiation was detected by an external bolometer. The signal in this latter configuration corresponds to a double-pass transmission. For the absolute transmission measurement in Fig. 2, the sample was mounted on a rotation stage that enabled referencing by an open aperture.
@@
Time-domain THz spectroscopy: Experiments using this method were carried out in reflection and transmission geometries. Reflection measurements, performed on sample B, were performed at an incidence angle of 11, using a THz optical setup with four off-axis parabolic mirrors. We used a photoconductive source and detector antennae, controlled with a standard THz time-domain spectrometer (Toptica). The sample was placed in an optical cryostat with quartz windows, and reflection spectra were collected as a function of the sample temperature in the range of 10-300 K. We calculated temperature-differential reflection spectra that show the $\alpha$ magnon mode because its frequency softens with increasing temperature. These spectra reveal strong interactions of the magnon with the modes of a Fabry-Perot cavity, formed due to the parallel-plane shape of the sample [50]. Due to a relatively large sample B thickness (989 $\mu$m), the cavity modes are dense in the frequency domain.
@@
The time-domain THz experiments in the transmission geometry were performed in two custom-made THz setups. For measurements without magnetic field, broadband pulses (0.2-3 THz) were generated by a photoconducting antenna and detected by electro-optic sampling in a ZnTe crystal. The sample was cooled in an optical cryostat (Optistat, Oxford instruments) from 300 down to 7 K. For more details on the setup, see Ref. [51]. For measurements in a magnetic field, broadband terahertz pulses were generated and detected using commercial fiber-coupled photoconductive switches operating with a femtosecond optical fiber laser system (TeraSmart, Menlo Systems GmbH). The sample was placed in an Oxford Instruments Spectromag He bath cryostat fitted with mylar windows and a superconducting coil, allowing for applying a magnetic field of up to 7 T, and cooling the sample down to 2 K. The measurements were performed in the Faraday geometry.
@@
THz magnetic-circular dichroism (THz-MCD): Temperature and magnetic-field-dependent transmittance measurements with a circularly polarized THz beam were performed in the Faraday geometry. The optically pumped THz/FIR gas laser (Edinburgh Instruments FIRL100) produced a highly monochromatic, linearly polarized THz beam, which was converted into a left-handed or right-handed circularly polarized state using a broadband tunable terahertz retarder [52]. The cyclotron resonance absorption of a high-quality two-dimensional electron gas in the GaAs quantum well is measured to ensure pure circular polarization. The magneto-optical cryogenic system (Oxford Instruments Spectromag SM4000) provides temperatures from 2 to 300 K and magnetic fields up to 10 T in both polarities. The relative transmittance is given as a ratio of the signal transmitted through the sample and detected by a liquid-helium-cooled bolometer to the signal monitoring the laser output. An additional set of data (see the inset of Fig. 4d) was collected using the described setup, but with linearly polarized radiation instead, to illustrate the fine tuning of the $\alpha$ magnon energy with $T$.
@@
Appendix E Modeling of time-domain THz data

### v2

@@ IV Results and discussion
Let us now compare the $B$-field dependence of the $\alpha$ mode energy with the proposed theoretical model. The observed blueshift matches well the expectations raised by Eq. 3. The validity of the theoretical model can be further corroborated using an additional set of data collected at various angles $\theta_{B}$ between $\mathbf{B}$ and the $c$ axis. Theoretically, we expect a weakening of the blueshift with increasing $\theta_{B}$, to gradually approach Eq. 5 with a nearly vanishing $B$-dependence for $E_{\mathrm{an}}\ll E_{J}$. Experimentally, we have indeed found that the blueshift of the $\alpha$ mode gradually flattens with increasing $\theta_{B}$. This is illustrated in Figs. 5a-e where magneto-transmission data collected in the Faraday configuration are presented, this time at varying angle between the applied magnetic field and $c$ axis. The blueshift of the $\alpha$ mode becomes at the limit of our experimental resolution around $\theta_{B}=60^{\circ}$. At higher angles, no blueshift is observed, including the limiting case of $\theta_{B}=90^{\circ}$ which we probed in the Voigt configuration (Fig. 5).
Figure 5: Relative magneto-transmission, $T_{B}/T_{B=0}$ collected in the Faraday configuration on sample A at liquid helium temperature, plotted as color plots for several values of the angle between the magnetic field and the $c$ axis of $\alpha$-MnTe: $\theta_{B}=0,35,50,60,70$ and 90 in panels (a) to (f). The data in panel (a) is the same as in Fig. 4a, but this time normalized by the zero-field transmission. In this way, the dark and light colors correspond to areas of $B$-induced absorption and transmission, respectively. The upper and lower dashed lines are theoretically expected energies of the $\alpha$ and $\beta$ modes, respectively, calculated numerically using the model presented in Sec. II for $D=36.4$ $\mu$eV, $g=2.023$ and $E_{an}=91$ $\mu$eV.
@@
To account for this, we have fitted the experimentally observed blueshift, see the dashed line in Fig. 4c, using the formula $(E_{\alpha}^{2}+\xi^{2}\mu_{B}^{2}B^{2})^{1/2}$, cf. Eq. 3. The extracted parameter $\xi=(1.99\pm 0.03)$ then allows us to calculate an approximate effective relation between the magnetic anisotropy and the factor $g$: $D\approx(\sqrt{3}E_{\alpha}/15)\sqrt{1-\xi^{2}/g^{2}}$. It is also worth noting that the parameter $\xi$ corresponds to the lowest possible value of the $g$ factor consistent with our magneto-optical data. The SOC in Mn-based compounds is typically weak and the effective $g$ factor should not significantly deviate from 2. Setting the maximum value at $g=2.1$, we get the following upper limit for the out-of-plane magnetic anisotropy, $D<150$ $\mu$eV.
@@
Another independent estimate of magnetic anisotropy can be obtained thanks to the AFMR experiment in the SF configuration (${\mathbf{B}}\perp c$). There, the $\alpha$ mode is expected to redshift approximately by $2(g\mu_{B}BD)^{2}/E^{3}_{\alpha}$ as one can infer from Eq. 5. Since no such shift was observed experimentally, we conclude it to be smaller than our experimental resolution (0.5 cm-1 @ 16 T). In this way, we obtain the upper bound for the magnetic anisotropy: $D<250$ $\mu$eV. This estimate is less strict, but is consistent with the one based on our AFMR experiments in the EC configuration.
@@
Using exchange parameters available in the literature for $\alpha$-MnTe is another way to obtain an estimate of the magnetic anisotropy. Szuszkiewicz et al. [25] mapped the magnon dispersion using inelastic neutron scattering and compared their results with the dispersion obtained by the LSW theory, considering a spin Hamiltonian equivalent to Eq. (1). They found good agreement with the theory, except for small departures around zone boundaries, and gave solid estimates of the corresponding exchange coupling constants. However, magnetic anisotropy seems to be the least precise parameter deduced, mainly due to the limited data set around $k=0$.
Taking the exchange parameters $J_{1}$ and $J_{3}$ from Szuszkiewicz et al. [25] we obtain $E_{J}\approx 33$ meV which is essentially the magnon bandwidth seen in Fig. 5 of that reference. Practically the same value ($E_{J}\approx 34$ meV) is also found in recent inelastic neutron scattering studies of $\alpha$MnTe by Liu et al. [13]. These effective exchange energies determined experimentally are not far from the value of 27 meV obtained numerically, within the density functional theory (more specifically, DFT+$U$) framework [2].
Using the experimentally determined $E_{J}$ [25, 13] and the energy of the $\alpha$ mode energy deduced from the optical measurements in this work (see Fig. 2), we conclude that the out-of-plane component of the magnetic anisotropy falls within the interval of $D=E_{\alpha}^{2}/(4SE_{J})=(40\pm 10)$ $\mu$eV, i.e., $E_{\mathrm{an}}=(100\pm 25)$ $\mu$eV. The declared error bar mainly reflects the uncertainty in the exchange parameters taken from the literature. Having estimated $D$, we may, in turn, evaluate the effective $g$ factor, $g=(2.03\pm 0.03)$ based on the AFMR data collected in the EC configuration, see the discussion above.
The deduced $D$ may be compared to both experimental and theoretical values determined in the previous work. In inelastic neutron scattering experiments on $\alpha$-MnTe, see Refs. [25] and [13], the mutually inconsistent values $(22\pm 3)$ $\mu$eV and $(48.3\pm 0.5)$ $\mu$eV, respectively, were reported. This suggests that the error bars of microscopic constants deduced in these experiments might be underestimated. A theoretical estimate of the single-ion magnetic anisotropy, $D_{\mathrm{theor}}\approx 20$ $\mu$eV can be found in Ref. [5]. It is also interesting to note that the absolute value of $D$ is comparable to the single-ion magnetic anisotropy in $\alpha$-MnTe lightly doped with lithium. However, the sign seems to be reversed, thus making Li-doped $\alpha$-MnTe easy-axis rather than easy-plane antiferromagnet [49, 50].
Knowing the magnetic anisotropy $D$ and the effective $g$ factor experimentally, we may test the observed magneto-optical response for various orientations of the magnetic field, see color plots in Fig. 5, against theoretical expectations. Although the theoretical model in Sect. II only provides us with simple analytical formulae (Eqs. 2-5) for special cases, $\theta_{B}=0$ and $\pi/2$, the numerical solution can be found for any $\theta_{B}$. The theoretically expected energies of both $\alpha$ and $\beta$ modes have been plotted as dashed lines in Fig. 5. The theoretical model compares well with our experimental data. Indeed, the blueshift of the $\alpha$-mode becomes below our experimental resolution for angles $\theta_{B}>60$ deg.
Let us close the discussion with another experimentally probed characteristic of the observed $\alpha$ mode, its partial circular polarization at $B>0$, see Fig. 6. It was observed using the THz MCD technique, which operates at a fixed laser frequency with a well-defined circular polarization, and follows the intensity of the transmitted radiation as a function of $B$. The partial circular polarization of the $\alpha$ mode is manifested by a profound difference in the observed absorption when the light propagates parallel or antiparallel to the direction of the magnetic field.
Magnetic circular dichroism in $\alpha$-MnTe is a complex phenomenon, sensitive to particular symmetries present (or broken), see, e.g. Ref. [7]. Nevertheless, the partial circular polarization of the $\alpha$ mode at $B>0$ could either be an effect of the $B$-induced magnetization or a consequence of the dynamics of magnetic moments (which move in elliptical rather than circular trajectories typical for easy-axis systems). In the ${\mathbf{B}}\|c$ configuration, the former appears due to spin canting away from the hexagonal plane (the EC regime). The angle between the spin and hexagonal planes reaches a few degrees in the highest applied $B$, $\arcsin[E_{Z}/(2E_{J})]$ (6), for the expected strength of the exchange coupling [25]. Note that the experiment with circularly polarized radiation was performed at higher temperatures ($T=200$ K) at which it was possible to achieve resonance between the field-dependent $\alpha$ mode and one of the (discrete) emission lines of the gas laser. This temperature-driven tuning of the magnon energy, at several values of the magnetic field, is demonstrated in the inset of Fig. 6.
Figure 6: Magneto-transmission of sample A, probed by the THz/FIR laser. The magnetic-field sweep ($\mathbf{B}\|c$) at a temperature of $T=200$ K was obtained using circularly polarized radiation with a photon energy of $\hbar\omega=2.87$ meV. The temperature sweeps, shown in the inset, were measured at several magnetic fields using a linearly polarized laser beam with a photon energy of 2.71 meV.
@@ V Conclusions
A long-wavelength ($k=0$) magnon mode was identified in the THz response of bulk $\alpha$-MnTe using the experimentally observed dependence of its energy on the applied magnetic field and temperature. The observed behavior aligns well with predictions from a simplified model of an easy-plane (hard-axis) antiferromagnet, developed within the framework of linear spin-wave theory. Its energy at low temperatures, 3.5 meV, ranks $\alpha$-MnTe among antiferromagnetic materials with moderate magnon gap (smaller than in some uniaxial systems [51, 52] but still appreciable). On the other hand, the magnon remains observable even at room temperature, which is close to $T_{N}$ of hexagonal MnTe. Combining our experimental results with the values of the exchange parameters published earlier, we determined the out-of-plane component of the magnetic anisotropy as $D=(40\pm 10)$ $\mu$eV and the effective $g$ factor as $g=(2.03\pm 0.03)$. These results are characteristic of bulk hexagonal MnTe and may serve as reference values for this material grown in the form of a thin film, where properties, such as the single-ion anisotropy, and consequently also magnetization, may be significantly impacted by the strain [53, 9].
@@ Acknowledgements.
We acknowledge discussions with M. Potemski and D. Pfannkuche. The work was supported by the Czech Science Foundation, project No. 23-04746S and TERAFIT project No. CZ.02.01.01/00/22__\__008/0004594 funded by OP JAK, call Excellent Research. Additional magnetometry on our bulk samples was performed by O. Kaman. J.D. and M.O. acknowledge the support received through the ANR-22-EXES-0001 (project CEQAS). Authors acknowledge the support of the LNCMI-CNRS, a member of the European Magnetic Field Laboratory (EMFL). Supplemental experiments were carried out in MGML (mgml.eu), supported by the Ministry of Education, Youth and Sports, the Czech Republic within the program of the Czech Research Infrastructures (project No. LM2023065). P.K. and L.N. acknowledge funding from the Grant Agency of the Charles University (grants No. 166123 and SVV--2024--260720). M.B. acknowledges the funding from the SONATA BIS-13 2023/50/E/ST3/00584 grant of the National Science Centre of Poland.
@@ Appendix A Linear spin wave theory
The classical ground state configuration has to be determined first. Writing the classical energy ${\cal E}$ associated with the Hamiltonian (1) in the macrospin representation we get for $\mathcal{E}/S$:
@@
where $\mathbf{m}_{i}$ denotes unit vectors corresponding to the respective sublattice magnetizations. The external magnetic field is $\mathbf{B}=\mathbf{b}B$, and without loss of generality, the unit vector $\mathbf{b}$ is taken to be in the plane $\{\mathbf{x},\mathbf{z}\}$. The quantities $E_{\mathrm{an}}$, $E_{Z}$ and $E_{J}$ are defined in the main text.
@@ Appendix B Temperature dependence of electrical resistivity
The temperature dependence of the electrical resistivity for samples A and B is plotted in Fig. 7. Qualitatively, both samples show the same behaviour with increasing $T$. A sharp decrease in resistance, with the minima below 100 K, followed by a gradual increase at higher temperatures. At high temperatures, the resistance of sample B appears to be approximately one order of magnitude higher than that of sample A. Assuming that the scattering of free carriers at high temperatures is dominated by intrinsic mechanisms (optical phonons), the difference in the resistance values may suggest a higher carrier density in sample A.
In terms of experimental details, the resistivity of sample A was measured using a PPMS system, with a maximum current of 100 $\mu$A, using low-frequency ac technique (4.5 Hz), using four contact made by silver epoxy. The resistivity of sample B was measured in a He cooled cryostat with a variable temperature inset. Alternating current of 10 $\mathrm{\mu}$A and frequency of 14.1 Hz was applied to a millimeter-size bar and longitudinal voltage drops were detected with lock-in voltmeters.
@@
Figure 7: Resistivity dependence on temperature for samples A and B with carrier freeze-out evidenced at low temperatures. The experiments were carried out using the conventional lock-in technique at low frequencies. The profile of $R(T)$ curve qualitatively match overall transmission of $\alpha$-MnTe samples in the THz range, as discussed in the main text.
Appendix C Description of experimental techniques
@@
Fourier transform magneto-spectroscopy: Nonpolarized radiation from a mercury lamp was analyzed by the Bruker Vertex 80v spectrometer and delivered, via light-pipe optics, to the sample which was kept at $T=4.2$ K in the helium exchange gas and placed inside a cryostat in a superconducting or resistive coil, below and above 16 T, respectively. In the superconducting coil, the radiation was then detected by a composite bolometer placed just below the sample. In this case, both Faraday and Voigt configurations were possible, with the light propagating along or perpendicular to the applied magnetic field, respectively. In the resistive coil, measurements were made only in the Faraday configuration. In order to reduce the noise level induced by the cooling water, the sample was placed on a mirror and the radiation was detected by an external bolometer. The signal in this latter configuration corresponds to a double-pass transmission. For the absolute transmission measurement in Fig. 2, the sample was mounted on a rotation stage that enabled reference using an open aperture.
@@
Time-domain THz spectroscopy: Experiments using this method were carried out in reflection and transmission geometries. Reflection measurements of sample B were performed at an incidence angle of about 9, using a THz optical setup with four off-axis parabolic mirrors. We used a photoconductive source and detector antennae, controlled with a standard THz time-domain spectrometer (Toptica). The sample was placed in an optical cryostat with quartz windows, and reflection spectra were collected as a function of the sample temperature in the range of 10-300 K. We calculated temperature-differential reflection spectra that show the $\alpha$ magnon mode because its frequency softens with increasing temperature. These spectra reveal strong interactions of the magnon with the modes of a Fabry-Perot cavity, formed due to the parallel-plane shape of the sample [54]. Due to a relatively large sample B thickness (989 $\mu$m), the cavity modes are dense in the frequency domain.
@@
The time-domain THz experiments in the transmission geometry were performed in two custom-made THz setups. For measurements without magnetic field, broadband pulses (0.2-3 THz) were generated by a photoconducting antenna and detected by electro-optic sampling in a ZnTe crystal. The sample was cooled in an optical cryostat (Optistat, Oxford Instruments) from 300 down to 7 K. For more details on the setup, see Ref. [55]. For measurements in a magnetic field, broadband terahertz pulses were generated and detected using commercial fiber-coupled photoconductive switches operating with a femtosecond optical fiber laser system (TeraSmart, Menlo Systems GmbH). The sample was placed in an Oxford Instruments Spectromag He bath cryostat fitted with mylar windows and a superconducting coil, allowing for the application of a magnetic field of up to 7 T, and cooling the sample down to 2 K. The measurements were performed in the Faraday geometry.
@@
THz magnetic-circular dichroism (THz-MCD): Temperature and magnetic-field-dependent transmittance measurements with a circularly polarized THz beam were performed in the Faraday geometry. The optically pumped THz/FIR gas laser (Edinburgh Instruments FIRL100) produced a highly monochromatic, linearly polarized THz beam, which was converted into a left-handed or right-handed circularly polarized state using a broadband tunable terahertz retarder [56]. The cyclotron resonance absorption of a high-quality two-dimensional electron gas in the GaAs quantum well is measured to ensure pure circular polarization. The magneto-optical cryogenic system (Oxford Instruments Spectromag SM4000) provides temperatures from 2 to 300 K and magnetic fields up to 10 T in both polarities. The relative transmittance is given as a ratio of the signal transmitted through the sample and detected by a liquid-helium-cooled bolometer to the signal monitoring the laser output. An additional set of data (see the inset of Fig. 6) was collected using the described setup, but with linearly polarized radiation instead, to illustrate the fine tuning of the $\alpha$ magnon energy with $T$.
@@
Appendix D Modeling of time-domain THz data